OpenStudy (df001):

Use the quadratic formula to find the complex roots (Leave answer in simplest a+bi form) : 3x²-2x+1=0

OpenStudy (anonymous):

deja vu that strange feeling you get when you think you've been there before

OpenStudy (anonymous):

didn't we just do this and get \[\frac{1\pm\sqrt{2}i}{2}\]

OpenStudy (df001):

Sorry, I had the question wrong

OpenStudy (anonymous):

if you want to make it look like \(a+bi\) split it up as \[\frac{1}{2}+\frac{\sqrt{2}}{2}i\]

OpenStudy (df001):

is that the simplest a+bi form?

OpenStudy (df001):

oh

OpenStudy (anonymous):

yes

OpenStudy (df001):

but, I was curious why 2 isnt multiplied by 3 on the denominator

OpenStudy (df001):

\[\frac{ 2\pm \sqrt{-8} }{ 2(3) }\]

OpenStudy (wolf1728):

a=3 b=-2 c=1 x=[-b +- sqrt(b^2 -4*a*c)] / 2*a x= [--2 +- sqrt(4 - 4*3*1)]/ 6 x = [2 +- sqrt(4 -12)]/6 x = [2 +- sqrt(8)] / 6

OpenStudy (df001):

can I convert to a+bi form

OpenStudy (df001):



OpenStudy (df001):